Find two consecutive integers such that the sum of their squares is 5. How do I solve this problem?
Answer by ptaylor(2198)

Let x=one integer
Then x+1 is the other integer
Now we are told that x^2+(x+1)^2=5 get rid of parens
x^2+x^2+2x+1=5 subtract 5 from both sides
x^2+x^2+2x+1-5=0 combine like terms
2x^2+2x-4=0 quadratic in standard form and it can be factored:
(2x+4)(x-1)=0
2x=-4
x=-2
x+1=-1
and
x=1
x+1=2
CK
(-2)^2+(-1)^2=5
4+1=5
5=5
and1^2+2^2=5
1+4=5
5=5
Hope this helps----ptaylor